OpenStudy (anonymous):

b. For the first phase of the project, imagine you want to cover the backyard with decorative rock and plant some trees. You need 30 tons of rock to cover the area. If each ton costs $60 and each tree is $84, what is the maximum number of trees you can buy with a budget of $2,500? Write an inequality that illustrates the problem and solve. Express your answer as an inequality and explain how you arrived at your answer.

OpenStudy (anonymous):

I know that 30 tons of rock is 1800, and after that 8 trees can be bought. I can't figure out how to write the equation.

OpenStudy (anonymous):

\[84t \le 700\] 84 per tree, and you have 700 left to spend \[t \le \frac{25}{3}\] so you can buy 8 full trees

OpenStudy (anonymous):

full equation is : \[30*60+84t \le 2500\]